40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 16 23:25:07 2017, gautam, Update, SUS, MC1 SUS electronics investigation 20170116_231625.pngIMG_7175.JPGIMG_7174.JPG
    Reply  Tue Jan 17 21:29:52 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
       Reply  Wed Jan 18 11:40:54 2017, gautam, Update, SUS, MC1 SUS electronics investigation MC1glitches.pngMC1coilGlitches.pngMC1CoilGlitchZoomed.png
          Reply  Wed Jan 18 14:23:47 2017, gautam, Update, SUS, MC1 SUS electronics investigation 
             Reply  Wed Jan 18 18:44:53 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCsatBoxSwitch.png
                Reply  Thu Jan 19 08:25:12 2017, Steve, Update, SUS, MC1 SUS electronics investigation MC1_MC3_ITMY_ETMX_sensors.pngsensors_UL.png
                   Reply  Thu Jan 19 12:00:10 2017, gautam, Update, SUS, MC1 SUS electronics investigation SatBoxSwitchGlitch.png
                      Reply  Thu Jan 19 19:56:09 2017, rana, Update, SUS, MC1 SUS electronics investigation 
                         Reply  Fri Jan 20 11:16:30 2017, gautam, Update, SUS, MC1 SUS electronics investigation MCglitchesReturn.png
                            Reply  Mon Jan 23 10:24:01 2017, Steve, Update, SUS, MC1 SUS electronics investigation 3glitchingSUS.png
Message ID: 12728     Entry time: Tue Jan 17 21:29:52 2017     In reply to: 12725     Reply to this: 12731
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 SUS electronics investigation 

 

Quote:
 

After some investigation, Rana found that on the AA filter board end, one of the 4pin LEMOs from the whitening board had one of its wires come unstuck from where it was soldered (this presumably happened while we were squishing cables tonight, as the LR channel was fine before that). Also, there was no heat shrink used on any of the solder joints.

The faulty cable has been re-soldered (with heat shrink) and replaced. All 5 sensor signals appear normal on dataviewer now. I am leaving things in this state for the night, let us see if the glitches return overnight.

PSL shutter remains closed

ELOG V3.1.3-